# PPNet: A Novel Neural Network Structure for End-to-End Near-Optimal Path Planning

### v1

@@ V Experiments
For training, we adopt a polynomial learning rate decay schedule and employ SGD as the optimizer. Momentum and weight decay are set to 0.9 and 0 respectively for all the experiments on the three datasets. We set an initial learning rate of 0.08 for all experiments. We set the batch size to 8 and the total iteration to 80,000 for experiments of comparison of different encoders of SegNet. And we used mmsegmentation [12] for building SegNet and experiments in Section V-A. The path planning problems used in this section are generated by EDaGe-PP, which contains up to 50 obstacles in each problem.
V-A Comparison Of Different Structures of SegNet
In this section, we test the performance of different structures of the segmentation model in path planning. For evaluating different structures of SegNet, the size of the dataset used in this section is 320,000.
@@ V-A1 Comparison Of Encoders
In the experiments on Table I, the reason that leads the performance of models with these encoders in path planning is quite different from the performance in classical CV tasks is the different attention mechanisms used in these encoders introduced in different assumptions, which have quite different influence in path planning.
@@
NAT uses Neighborhood Attention (NA) [15]. NA can be seen as a sliding window variant of SWSA. In NA, each patch only calculates attention with the adjacent patches which makes the computation more efficient, and the correlation of adjacent patches tend to be higher than SA. Such an assumption is reasonable for path planning. And in our experiment, the model using NAT as an encoder obtained the best performance.
DiNAT uses Dilated NA (DiNA) [16] which is inspired by dilated convolution that can make the window of NAT extremely large with the same amount of computation. With the larger window, the model can obtain the receptive field as large as the task need. But in DiNA, patches can not calculate attention directly with adjoining patches after the first layer which makes the model using DiNAT can not converge in path planning.
In summary, the model using NAT obtained the best performance due to the assumption which is introduced by NA. Nevertheless Swin outperform ViT in all classical task in CV, models using Swin and ViT perform similarly in path planning. Model using DiNAT can not converge in path planning.
@@
V-A2 Comparison Of decoders
Following the configuration of decoders above, the number of parameters of NAT-PUP-512, and NAT-UPerHead-512 are 104.1M, and 123.2M respectively. And the numbers of FLOPs of NAT-PUP-512, and NAT-UPerHead-512 are 53.47G, and 54.16G respectively. In the experiment on Table II, all NAT-UPerHead models outperform NAT-PUP-512. As the performance of NAT-UPerHead models is quite similar, we choose the NAT-UPerHead-64 as the model for SegNet.
@@ V-A2 Comparison Of decoders
PPNet-RRT* 25.6min (+14.2h) $\mathbf{37.3\pm 9.44}$ $40.80$
PPNet-IRRT* 10.1min (+14.1h) $37.8\pm 8.20$ $48.84$
PPNet-BIT* 5.96min (+16.3h) $37.8\pm 8.19$ $45.21$
PPNet-ABIT* 5.86min (+16.0h) $37.9\pm 8.04$ $40.92$
PPNet-EDaGe-PP 0.963min (+0.425h) $39.0\pm 7.39$ $\mathbf{94.47}$
@@ V-B EDaGe-PP Comparison with Popular Data Generation Methods
In the comparison of data generation methods, we divide the data generation time into two parts, which are the time for generating solutions and the time for generating data with the format for training of PPNet. For evaluating the time of generating solutions, we let RRT*, IRRT*, BIT*, and ABIT* run until they find a solution of Euclidean cost within the 5% range of the cost of the PPNet solution. As shown in Table III, the data generation methods that apply RRT*, IRRT*, BIT*, and ABIT* for finding solutions have about 27×27$\times$27 ×, 10×10$\times$10 ×, 6×6$\times$6 ×, 6×6$\times$6 × computation time for generating solutions respectively compared to EDaGe-PP, and have about $\times3$, $\times3$, 38×38$\times$38 ×, 38×38$\times$38 × computation time for generating data respectively compared to EDaGe-PP. For the cost of the solutions, We report mean computation times with standard deviation over 10,000 path planning problems. Due to the solutions which are generated by EDaGe-PP being continuous-curvature, PPNet can not output solutions comprised of broken lines, which makes the cost of the solutions that are found by PPNet-EDaGe-PP slightly higher than others. But because of the continuous-curvature solutions, the success rate of PPNet-EDaGe-PP is about 2×2$\times$2 × compared to the other methods.
@@ V-D PPNet Comparison With Learning-Based Planner
In this section, we evaluate PPNet against the representative learning-based planner MPNet. We choose MPNetPath:NP (B), called MPNetPath in this paper, which is the variant with the best performance of all variants of MPNet in computation time, for the experiments in this section. For following the settings of the dataset of MPNet, the size of the dataset used in this section is 4,000. The maximum number of iterations of MPNetPath is set to 130, and the part of replanning is 50.
Table IV present the numerical result of comparison of PPNet and the five variants of MPNetPath that are trained by the datasets generated by EDaGe-PP and the methods that apply classical planners, including RRT*, IRRT*, BIT*, and ABIT*, which are represented by MPNetPath-EDaGe-PP, MPNetPath-RRT*, MPNetPath-IRRT*, MPNetPath-BIT*, MPNetPath-ABIT* respectively. As shown in Table IV, PPNet outperforms all variants of MPNetPath in computation time, path cost, and success rate. And in all variants of MPNetPath, MPNetPath-PPNet outperforms other variants in path cost. But MPNetPath-EDaGe-PP obtains the worst performance in success rate. The reason is RNN that the neural networks used in MPNet have the limitation in predicting long sequence results, and the solutions generated by EDaGe-PP have more waypoints than other methods. But this feature is suitable for the methods that apply the neural networks with structures for CV, which is exhibited in the success rate of PPNet that is about 2×2$\times$2 × compared to MPNetPath. As for computation time, because of the end-to-end way that PPNet uses for solving path planning problems, PPNet provides about 7×7$\times$7 ×, 9×9$\times$9 ×, 9×9$\times$9 ×, 8×8$\times$8 ×, and 8×8$\times$8 × computation speed improvements compared to MPNetPath-RRT*, MPNetPath-IRRT*, MPNetPath-BIT*, MPNetPath-ABIT*, MPNetPath-EDaGe-PP respectively.
@@
V-E PPNet Comparison With Sampling-Based Planners
We let RRT*, IRRT*, BIT*, and ABIT* run until they find an equivalent solution to PPNet and a solution of Euclidean cost within the 2%, and 5% range of the cost of the PPNet solution respectively. We report mean computation times with standard deviation over 10,000 random path planning problems. PPNet can find a solution within a specific planning time, which is 15.3ms. The state-of-the-art classical planners, RRT*, IRRT*, BIT*, and ABIT* exhibit shorter computation times in simple planning problems, but higher mean computation times than PPNet. Furthermore, standard deviations of computation times are high, which is unacceptable in some tasks (Fig.5). As shown in Table V, RRT*, IRRT*, BIT*, and ABIT* need about 15×15$\times$15 ×, 8×8$\times$8 ×, 8×8$\times$8 ×, and 8×8$\times$8 × computation time for finding an equivalent solution to PPNet respectively. And Fig.5 presents a difficult case for sampling-based planners. For solving this path planning problem, RRT*, IRRT*, BIT*, and ABIT* need $2.452\mathrm{s}$, $0.580\mathrm{s}$, $60.001\mathrm{s}$, and $9.564\mathrm{s}$ of computation time. And PPNet can still output the near-optimal solution in 15.3ms.
VI Conclusion
In this article, we propose EDaGe-PP, which is an efficient data generation method for path planning. And due to the continuous-curvature path with analytical expression, the dataset is beneficial for the training of neural networks. We also propose PPNet, which is a two-level cascade neural network that each level solves one of the two subproblems of the path planning problem. PPNet can find a near-optimal solution in an end-to-end way within 15.3ms.
In future work, we will try to improve our neural network structure to develop an end-to-end near-optimal path planner in 3D space. In addition, extending EDaGe-PP to 3D space is also a challenging work.
Code is available at: https://github.com/AdamQLMeng/PPNet

### v2

@@ V Experiments
For training, we adopt a polynomial learning rate decay schedule and employ SGD as the optimizer. Momentum and weight decay are set to 0.9 and 0 respectively for all the experiments on the three datasets. We set an initial learning rate of 0.08 for all experiments. We set the batch size to 8 and the total iteration to 80,000 for experiments of comparison of different encoders of SpaceSegNet. And we used mmsegmentation [12] for building SpaceSegNet and experiments in Section V-A. The path planning problems used in this section are generated by EDaGe-PP, which contains up to 50 obstacles in each problem. In the experiments of previous learning-based methods, they test the methods only in a few maps, eg. NRRT* in 4 maps, and MPNet in 10 maps. Such a setting is reasonable for classic methods, but learning-based methods need sufficient different maps to validate the generalization. Therefore, we evaluate our methods on a dataset that each data is a path planning problem in a different map, which can validate the generalization of our method. Unlike the previous works combining machine learning techniques and sampling-based methods, our method is a deep learning method without implementing traditional path planning methods, and our method can only solve path planning problems in 2D scenarios. Developing methods for solving path planning problems in higher dimension space in an end-to-end way will be our future work.
V-A Comparison of Different Structures of SpaceSegNet
In this section, we test the performance of different structures of the segmentation model in path planning. For evaluating different structures of SpaceSegNet, the size of the dataset used in this section is 320,000, which means the dataset contains 320,000 different maps. Unlike the traditional segmentation tasks that can segment target regions by texture in images, eg. ADE20K, and crack segmentation, the space of the target path should be completely in the free space, which means the target region is in the white region has no texture, this makes the task much more difficult.
@@ V-A1 Comparison of Encoders
In the experiments on Table I, the reason that the performance of models with these encoders in path planning is quite different from the performance in classical CV tasks is the different attention mechanisms used in these encoders introduced in different assumptions, which have quite different influence in path planning.
@@
NAT uses Neighborhood Attention (NA) [15]. NA can be seen as a sliding window variant of SWSA. In NA, each patch only calculates attention with the adjacent patches which makes the computation more efficient, and the correlation of adjacent patches tends to be higher than SA. Such an assumption is reasonable for path planning. And in our experiment, the model using NAT as an encoder obtained the best performance.
DiNAT uses Dilated NA (DiNA) [16] which is inspired by dilated convolution that can make the window of NAT extremely large with the same amount of computation. With the larger window, the model can obtain the receptive field as large as the task needs. But in DiNA, patches can not calculate attention directly with adjoining patches after the first layer which makes the model using DiNAT can not converge in path planning.
In summary, the model using NAT obtained the best performance due to the assumption which is introduced by NA. Nevertheless Swin outperforms ViT in all classical tasks in CV, and models using Swin and ViT perform similarly in path planning. Model using DiNAT can not converge in path planning.
@@
V-A2 Comparison of Decoders
Following the configuration of decoders above, the number of parameters of NAT-PUP-512, and NAT-UPerHead-512 are 104.1M, and 123.2M respectively. The numbers of FLOPs of NAT-PUP-512, and NAT-UPerHead-512 are 53.47G, and 54.16G respectively. In the experiment on Table II, all NAT-UPerHead models outperform NAT-PUP-512. As the performance of NAT-UPerHead models is quite similar, we choose the NAT-UPerHead-64 as the model for SpaceSegNet.
@@ V-A2 Comparison of Decoders
RRT* 25.6min (+14.2h) $\mathbf{37.3\pm 9.44}$ $40.80$
IRRT* 10.1min (+14.1h) $37.8\pm 8.20$ $48.84$
BIT* 5.96min (+16.3h) $37.8\pm 8.19$ $45.21$
ABIT* 5.86min (+16.0h) $37.9\pm 8.04$ $40.92$
EDaGe-PP 0.963min (+0.425h) $39.0\pm 7.39$ $\mathbf{94.47}$
@@ V-B EDaGe-PP Comparison with Popular Data Generation Methods
In the comparison of data generation methods, we divide the data generation time into two parts, which are the time for generating solutions and the time for generating data with the format for training of PPNet. For evaluating the time of generating solutions, we let RRT*, IRRT*, BIT*, and ABIT* run until they find a solution of Euclidean cost within the 5% range of the cost of the PPNet solution. As shown in Table III, the data generation methods that apply RRT*, IRRT*, BIT*, and ABIT* for finding solutions have about 27×27$\times$27 ×, 10×10$\times$10 ×, 6×6$\times$6 ×, 6×6$\times$6 × computation time for generating solutions respectively compared to EDaGe-PP, and have about $\times3$, $\times3$, 38×38$\times$38 ×, 38×38$\times$38 × computation time for generating data respectively compared to EDaGe-PP. For the cost of the solutions, We report mean computation times with standard deviation over 10,000 path planning problems. Due to the solutions that are generated by EDaGe-PP being continuous-curvature, PPNet can not output solutions comprised of broken lines, which makes the cost of the solutions that are found by PPNet-EDaGe-PP slightly higher than others. However, because of the continuous-curvature solutions, the success rate of PPNet-EDaGe-PP is about 2×2$\times$2 × compared to the other methods.
@@ V-D PPNet Comparison With Learning-Based Planner
In this section, we evaluate PPNet against the representative learning-based planner MPNet. We choose MPNetPath:NP (B), called MPNetPath in this paper, which is the variant with the best performance of all variants of MPNet in computation time, for the experiments in this section. For following the settings of the dataset of MPNet, the size of the dataset used in this section is 4,000, but the dataset contains 4,000 different maps, which is much larger than the setting in the experiments of MPNet (number of maps: 10). The maximum number of iterations of MPNetPath is set to 130, and the part of replanning is 50.
Table IV present the numerical result of comparison of PPNet and the five variants of MPNetPath that are trained by the datasets generated by EDaGe-PP and the methods that apply classical planners, including RRT*, IRRT*, BIT*, and ABIT*, which are represented by MPNetPath-EDaGe-PP, MPNetPath-RRT*, MPNetPath-IRRT*, MPNetPath-BIT*, MPNetPath-ABIT* respectively. As shown in Table IV, PPNet outperforms all variants of MPNetPath in computation time, path cost, and success rate. In all variants of MPNetPath, MPNetPath-PPNet outperforms other variants in path cost. But MPNetPath-EDaGe-PP obtains the worst performance in success rate. The reason is RNN that the neural networks used in MPNet have the limitation in predicting long sequence results, and the solutions generated by EDaGe-PP have more waypoints than other methods. However, this feature is suitable for the methods that apply the neural networks with structures for CV, which is exhibited in the success rate of PPNet which is about 2×2$\times$2 × compared to MPNetPath. As for computation time, because of the end-to-end way that PPNet uses for solving path planning problems, PPNet provides about 7×7$\times$7 ×, 9×9$\times$9 ×, 9×9$\times$9 ×, 8×8$\times$8 ×, and 8×8$\times$8 × computation speed improvements compared to MPNetPath-RRT*, MPNetPath-IRRT*, MPNetPath-BIT*, MPNetPath-ABIT*, MPNetPath-EDaGe-PP respectively.
V-E PPNet Comparison With Sampling-Based Planners
We let RRT*, IRRT*, BIT*, and ABIT* run until they find an equivalent solution to PPNet and a solution of Euclidean cost within the 2%, and 5% range of the cost of the PPNet solution respectively. We report mean computation times with a standard deviation over 10,000 random path planning problems. PPNet can find a solution within a specific planning time, which is 15.3ms. The state-of-the-art classical planners, RRT*, IRRT*, BIT*, and ABIT* exhibit shorter computation times in simple planning problems, but higher mean computation times than PPNet. Furthermore, standard deviations of computation times are high, which is unacceptable in some tasks (Fig.5). As shown in Table V, RRT*, IRRT*, BIT*, and ABIT* need about 15×15$\times$15 ×, 8×8$\times$8 ×, 8×8$\times$8 ×, and 8×8$\times$8 × computation time for finding an equivalent solution to PPNet respectively. And Fig.5 presents a difficult case for sampling-based planners. For solving this path planning problem, RRT*, IRRT*, BIT*, and ABIT* need $2.452\mathrm{s}$, $0.580\mathrm{s}$, $60.001\mathrm{s}$, and $9.564\mathrm{s}$ of computation time. And PPNet can still output the near-optimal solution in 15.3ms.
VI Conclusion
In this article, we also propose PPNet, which is a two-stage neural network in which each stage solves one of the two subproblems of the path planning problem. Moreover, we also propose EDaGe-PP, an efficient data generation method for path planning. Due to the continuous-curvature path with analytical expression, the dataset is beneficial for the training of neural networks. PPNet can find a near-optimal solution in an end-to-end way within 15.3ms.
In future work, we will try to improve our neural network structure to develop an end-to-end near-optimal path planner in 3D space. In addition, extending EDaGe-PP to 3D space is also challenging work. Code is available at: https://github.com/AdamQLMeng/PPNet.
@@ VI Conclusion
Figure 6: An example of PPNet, RRT*, IRRT*, BIT*, and ABIT* run on a u-turn environment. Each algorithm was run until it found an equivalent solution to PPNet ($c=50.03$). (a) $t=0.015\mathrm{s}$. (b) $t=0.539\mathrm{s}$. (c) $t=0.503\mathrm{s}$. (d) $t=0.220\mathrm{s}$. (e) $t=0.104\mathrm{s}$.